# C-MELT: Contrastive Enhanced Masked Auto-Encoders for ECG-Language Pre-Training

### v2

@@ 4 Experiments
Table 1: Performance for 5 lead combinations in diagnosis classification (Dx.) and patient identification (Id.). P-N-lead indicates N zero-padded unavailable leads. CinC scores (Dx.) scaled by 100.
@@
4.1 Implementation details
4.1.1 Pre-training Task.
Pre-train Dataset.
Experimental Configurations.
Our proposed model is developed based on the fairseq-signals framework in our work. We select the Adam optimizer with a learning rate of $5\times 10^{-5}$ and use a tri-stage scheduler with ratios of 0.1, 0.4, and 0.5 for learning rate adjustments. The optimizer is configured with $\beta_{1}=0.9$, $\beta_{2}=0.98$, an epsilon value of $1\times 10^{-6}$, and a weight decay of 0.01. We pre-train the proposed model for 300000 steps, maintaining a batch size of 128. The quantitative experiments are conducted on a single NVIDIA H100-80GB GPU.
@@ 4.1 Implementation details
Table 2: Performance comparison across multiple methods and datasets. The results are shown for different percentages of training data used (1%, 10%, 100%).
Methods CSN PTBXL-Rhythm PTBXL-Form PTBXL-Sub PTBXL-Super CPSC2018
1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100%
SimCLR (Chen et al., 2020) 59.02 67.26 73.20 51.41 69.44 77.73 54.98 56.97 62.52 60.84 68.27 73.39 63.41 69.77 73.53 59.78 68.52 76.54
BYOL (Grill et al., 2020) 54.20 71.92 74.69 41.99 74.40 77.17 48.73 61.63 70.82 57.16 67.44 71.64 71.70 73.83 76.45 60.88 74.42 78.75
BarlowTwins (Zbontar et al., 2021) 60.72 71.64 77.43 50.12 73.54 77.62 52.12 60.39 66.14 62.57 70.84 74.34 72.87 75.96 78.41 55.12 72.75 78.39
MoCo-v3 (Chen et al., 2021) 54.61 74.26 77.68 51.38 71.66 74.33 50.32 63.71 71.31 55.88 69.21 76.69 73.19 76.65 78.26 62.13 76.74 75.29
SimSiam (Chen & He, 2021) 58.25 68.61 77.41 49.30 69.47 75.92 55.16 62.91 71.31 62.52 69.31 76.38 73.15 72.70 75.63 58.35 72.89 75.31
TS-TCC (Eldele et al., 2021) 55.26 68.48 76.79 43.34 69.48 78.23 48.04 61.79 71.18 53.54 66.98 77.87 70.73 75.88 78.91 57.07 73.62 78.72
CLOCS (Kiyasseh et al., 2021) 54.38 71.93 76.13 47.19 71.88 76.31 51.97 57.79 72.65 57.94 72.55 76.24 68.94 73.36 76.31 59.59 77.78 77.49
ASTCL (Wang et al., 2023) 56.40 70.87 75.79 52.38 71.98 76.05 44.14 60.93 66.99 61.86 68.77 76.51 72.51 77.31 81.02 57.90 77.01 79.51
CRT (Zhang et al., 2023) 56.21 73.70 78.80 47.44 73.52 74.41 46.41 59.49 68.73 61.98 70.82 78.67 69.68 78.24 77.24 58.01 76.43 82.03
ST-MEM (Na et al., 2024) 59.77 66.87 71.36 51.12 65.44 74.85 55.71 59.99 66.07 54.12 57.86 63.59 61.12 66.87 71.36 56.69 63.32 70.39
MERL (Liu et al., 2024) 66.60 82.74 87.95 53.33 82.88 88.34 58.26 72.43 79.65 64.90 80.56 84.72 82.39 86.27 88.67 70.33 85.32 90.57
C-MELT 80.04 87.36 90.71 86.61 92.83 96.71 70.10 78.91 83.98 77.74 82.92 85.15 83.15 88.36 90.11 85.46 91.35 94.92
@@
PhysioNet 2021. This dataset contains ECG samples (500 Hz) ranging between 5 and 144 seconds. We process and fine-tune the subsets as described in (Oh et al., 2022) to validate the pre-trained ECG encoder in two downstream tasks: 1) 26-multi-label cardiac arrhythmia classification (Dx.); 2) patient identification (Id.), predicting patient ownership of ECG recordings.
PTB-XL. The PTB-XL dataset includes 21,837 ECG signals collected from 18,885 patients. Each sample has a 12-lead ECG recording sampled at 500 Hz over 10 seconds and corresponding cardiac labels. We follow (Liu et al., 2024) to split the dataset, including four sub-groups (super, sub, form, and rhythm). We consider them as the four separated datasets and prepare each of them with the same train, val, and test set split as in the original paper (Wagner et al., 2020).
CSN. This dataset consists of 23,026 ECG recordings sampled at 500 Hz for 10 seconds with 38 distinct labels. Therefore, it also supports the evaluation in a classification task. We use 70%:10%:20% data split as processed in (Liu et al., 2024).
CPSC2018. The dataset contains 6,877 standard 12-lead ECG recordings (500 Hz), which cover 9 distinct categories. Similarly, we also use the same data configuration following (Liu et al., 2024).
CODE-test: This contains 827 12-lead ECG samples (400 Hz) at varying lengths covering 6 abnormalities, annotated by several experienced residents and medical students. We resample the ECG signals to 500 Hz and adjust the lengths to 10 seconds.
Experimental Configurations.
To evaluate our model’s performance on downstream tasks, we conduct three experiments: 1) First, we integrate a linear layer on top of the pre-trained ECG encoder and fine-tune the entire model to test its efficacy in two tasks within the Physionet 2021 dataset: Dx. (by CinC score) and Id. (by % accuracy). We report the results with five cases of lead combinations, as presented in (Oh et al., 2022); 2) Second, we also implement a linear classifier but keep the ECG encoder frozen. This linear probing approach is applied at different training set sizes (1%, 10%, and 100%) to assess the macro AUC score (%) on the PTB-XL, CSN, and CPSC2018 test datasets, facilitating a comparison with our baseline (Liu et al., 2024); 3) Finally, we investigate zero-shot classification (AUC) on PTB-XL, CSN, CPSC2018 and CODE-test datasets. Here, the texts used are obtained by passing the category names through GPT-4o for capturing better medical context. The detailed configuration on each experiment is mentioned in Appendix A.1.
Table 3: Zero-shot performance comparison across multiple datasets.
@@
C-MELT 76.2 75.9 66.1 88.6 80.1 76.3 77.1
4.2 Quantitative Results
Full Fine-tuning Classifier.
As shown in Table 1, our method consistently outperforms previous approaches in both examined tasks. In the classification task, our model achieves 85.7% accuracy with all 12 leads, significantly higher than the best baseline (W2V+CMSC+RLM), which is 73.2%. This number is even lower than our setting with only 1 lead usage (76.5%). Interestingly, the 3-lead combination yields the second-highest result, only 1.5% lower than using all leads, while the 2-lead and 6-lead combinations produce comparable results, both around 81.5%. This suggests that the selected leads (I, II, V2) capture sufficient information for accurate performance. A similar pattern emerges in the identification task, where our model achieves 41.1% accuracy with a single lead, 60.5% with 3 leads, and 65.4% with 12 leads, surpassing the best baseline by 7%.
Linear Probing Classifier.
Zero-shot Classifier.
We first compare our method with MERL in conventional zero-shot settings across six datasets, as shown in Table 3. On average, our method achieves 77%, outperforming MERL by 2%. Notably, MERL performs particularly impressive on the CPSC2018 dataset, while its results on the other five datasets are consistently lower than ours. Next, we extend the comparison of our method with MERL and other SSL baselines under data distribution shifts. Specifically, we compare linear probing (100% training size) of SSL methods with MERL’s and our zero-shot approach. In this setup, source domain and target domain share some common categories. Details on this implementation can be found at Appendix A.1. As shown in Table 4, our results surpass MERL and other methods, except when CPSC2018 is the target domain, which aligns with our previous observations. Finally, Table 5 shows that our zero-shot model outperforms three experienced cardiologists (over 3%) and also the in-domain model (Ribeiro et al., 2020), i.e., trained with millions of annotated ECG examples. We will discuss more on zero-shot settings in the Appendix A.3.
Table 4: Zero-shot performance under data distribution shift.
@@ 4.2 Quantitative Results
MERL (Liu et al., 2024) 0% 88.21 78.01 76.77 76.56 74.15 82.86
C-MELT 0% 72.09 79.11 77.12 82.91 76.24 80.10
Table 5: ECG interpretation comparison: Human experts vs. DNN (Ribeiro et al., 2020) vs. C-MELT.
Cardio Resident Emergency Resident Medical Student DNN C-MELT (Zero-shot)
@@
4.3 Ablation Studies
We evaluate the impact of the key model components, the choice of language encoders, and varying the number of transformer layers in the ECG encoder for ablation studies. Here, we focus on three downstream tasks, including full fine-tuned diagnosis classification (results across five lead combinations), linear probing at 1% training size, and zero-shot classification with category names (results across PTB-XL, CSN, and CPSC2018 datasets).
Effects of Key Components.
To assess the contribution of different model components, including Flan-T5, Siglep, and N3S, we systematically remove one component at a time from the default proposed model. Specifically, we start by eliminating the N3S and train the model with randomly selected negative samples. Subsequently, we take the Siglep loss away to assess its effectiveness in capturing rich representative embeddings in both encoders. Lastly, by replacing the Flan-T5 language encoder with a standard Bert-base architecture (Devlin, 2018), we consider this as the baseline model. Table 7 demonstrates the results of this experiment. It can be seen that Siglep significantly enhances performance, showing an improvement of 15% over the base model. Meanwhile, adding N3S improves zero-shot classification by 2%, and introducing Flan-T5 enhances performance in linear probing by 4% compared to the baseline. These results underscore the effectiveness of each component in optimizing the model’s performance.
@@ 4.3 Ablation Studies
To better understand how our method improves downstream performance, we visualize and compare the t-SNE embeddings generated by our ECG encoder on the CSN test set with those from MERL. For clearer visualization, we include only samples from unique categories and exclude categories with fewer than 50 samples. Figure 3 reveals that our embeddings show more well-defined and distinct clusters representing different ECG diagnoses, which aligns with expectations.
Table 6: Effects of model components: ① FlanT5, ② Siglep, ③ N3S.
@@
Table 7: Effects of different language encoders.
Lang encoder Full fine-tune Linear probing Zero-shot
@@
Table 8: Effects of the number of transformer layers from ECG encoder. By default, our model contains 8 transformer layers.
@@
Figure 3: T-SNE visualization on the CSN test set.
Choice of Language Encoders.
Choice of Number ECG Transformer Layers.
As part of our ablation study, we explore the impact of varying the number of transformer layers (1, 4, 8) in the ECG encoder. As shown in Table 8, increasing the number of layers significantly improves performance. Specifically, the 1-layer model performs 11% worse than the 8-layer model in full fine-tuning and 13% worse in linear probing. For zero-shot, the 8-layer model still delivers superior results, with 2% and 3% higher performance than the 4-layer and 1-layer models, respectively. Although these differences are smaller than in full fine-tuning, they highlight the language encoder’s impact in improving performance.
@@ 5 Conclusion
We propose C-MELT to pre-train a model on ECG signals and corresponding texts, utilizing a novel contrastive masked transformer-based architecture. Our approach is generative self-supervised learning, enhanced with Siglep loss, and nearest-neighbor negative sampling to support contrastive aspects. Experimental results demonstrate that our method outperforms previous approaches across multiple datasets and on a range of downstream tasks, including under full fine-tuning, linear probing, and zero-shot classification. C-MELT shows promise in advancing ECG-based diagnostic models, paving the way for more accurate, efficient, and personalized cardiac care.
@@ Appendix A Appendix
In this section, we first visualize representative examples of ECG-text pairs from the MIMIC IV ECG dataset (Gow et al., 2023), as shown in Figure 4. We also indicate the top 30 common unique reports (before merging) in Figure 5. Prominent terms such as ”abnormal ecg”, ”normal ecg”, ”atrial fibrillation”, and ”sinus tachycardia” indicate common diagnoses, which suggests prevalent cardiovascular conditions and typical annotations within this dataset.
@@
Figure 5: WordCloud visualization on the top 30 common unique reports from MIMIC IV ECG dataset.
Next, we provide more details on data configurations in Table 9, including data split, number of classes, metrics, and the corresponding tasks with the given downstream dataset.
CODE-test:
Particularly, this data is from the work of Ribeiro et al. (2020), which is the test set used for evaluating their trained model’s performance compared with cardiology resident medical doctors. It is worth noting that their training set consists of over 2 million ECG records from 1,676,384 different patients in 811 counties. We evaluate the performance of our method on the same released test set of 827 samples in a zero-shot manner. These samples are originally sampled at 400 Hz, with durations of either 10 seconds or 7 seconds. Therefore, we resampled to 500 Hz and adjusted by truncating or padding with zeros as needed to get 10-second samples. For the gold standard (ground truth), two expert cardiologists provided their diagnoses. If they agree with each other, their consensus becomes the gold standard. In cases of disagreement, a third specialist reviews their diagnoses and determines the final decision.
@@
Table 10: Details on training configurations on the fine-tuned datasets. For optimizer, we keep using Adam in all experiments.
@@
For the distribution shift experiment, we follow the SCP-codes (classes) matching settings in (Liu et al., 2024), which can be seen in Table 11. This is to support three dataset matches (PTBXL-Super and CPSC2018), (PTBXL-Super and CSN), and (CPSC2018 and CSN). It is worth noting that the None value indicates the target dataset does not have a matching label for given labels in the source dataset.
@@
Why Using ETM Only Is Not A True Way To Zero-shot Learning.
As mentioned in the Method section, ETM functions as a contrastive learning technique in the masked auto-encoder architecture. However, it heavily relies on binary classification tasks with explicit ECG-text pairs to learn cross-modal correspondences. It is not designed for zero-shot learning which strongly requires the model to generalize to unseen tasks or classes without the need for such supervised pairings or fused information during training. This motivates us to use Siglep, boosting the model’s zero-shot ability.
@@
In medical datasets, particularly the MIMIC-IV ECG dataset (Gow et al., 2023), we observe a significant amount of duplicate or highly similar text samples: among nearly 800,000 records, only approximately 180,000 are unique. For instance, over 100,000 samples share an identical text report, which is ”sinus rhythm normal ecg”. Randomly selecting negative samples for contrastive loss training is not a suitable approach in this scenario. Therefore, we propose using the N3S technique to more effectively differentiate between similar and dissimilar samples, improving contrastive learning by selecting more meaningful negatives. Notably, we observe that during training, the ETM accuracy without N3S stagnates around 75% while with N3S, it exceeds 96%, demonstrating the significant impact of this approach.
A.3 Enhancing zero-shot performance with LLM.
Why Using LLMs But Not As MERL.
@@
In zero-shot learning, models typically rely on category names alone to make predictions. However, by incorporating Large Language Models (LLMs), we can enhance the context by generating richer, clinically relevant descriptions of the categories, as discussed in MERL (Liu et al., 2024). However, we observe two main drawbacks in their enhanced text reports, as shown in Figure 6: 1) MERL’s performance heavily depends on their sub-types and attributes searching prompt and additional database. This leads to a limitation when testing detailed analysis with labels that are different sub-types themselves. Moreover, this also raises suspicion about the performance when new tasks require labels that are not able to search sub-types and attributes in the database; 2) Following that point, MERL’s enhanced texts might be uncontrollable to the outputs where the LLMs provide wrong sub-types or unnecessary context. For example, ”Atrial Fibrillation” is already in ”AFIB” type but shown to misleadingly be in ”AF”-”Atrial Flutter”.
How Our Work Leverages LLM’s Strength.
We address these points using a straightforward prompt strategy with explicit instructions. Specifically, we employ a prompt: ”You are an experienced cardiologist. For a given clinical term such as ’normal ECG’, your job is to describe each term clinically and apply your medical domain knowledge to include other relevant explanations that will help a text encoder like Flan-T5 fully understand medical concepts. Do not include any recommendations in the description.” This makes the LLM generate clinically accurate and more focused explainable descriptions, enhancing the text encoding without introducing irrelevant or redundant information. For example, with the code ”AFIB”, our prompt on GPT-4o can output: ”Atrial Fibrillation (AFIB). Irregular and often rapid heart rate due to uncoordinated atrial activity.”.

### v3

@@ 4 Experiments
4.1 Implementation Details.
4.1.1 Pre-training Task.
Pre-train Dataset.
Experimental Configurations.
For model training, we use the Adam optimizer with a learning rate of $5\times 10^{-5}$ and use a tri-stage scheduler with ratios of 0.1, 0.4, and 0.5 for learning rate adjustments. The optimizer is configured with $\beta_{1}=0.9$, $\beta_{2}=0.98$, an epsilon value of $1\times 10^{-6}$, and a weight decay of 0.01. We pre-train the proposed model for 300000 steps, maintaining a batch size of 128. The quantitative experiments are conducted on a single NVIDIA H100-80GB GPU.
Table 1: Performance for 5 lead combinations in diagnosis classification (Dx., by CinC scores scaled by 100) and patient identification (Id., by %). P-N-lead indicates N zero-padded unavailable leads.
@@
Table 2: Performance comparison (AUC in %) across multiple methods and datasets. The results are shown for different percentages of training data used (1%, 10%, 100%).
Methods PTBXL-Super PTBXL-Sub PTBXL-Form PTBXL-Rhythm CPSC2018 CSN
1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100%
SimCLR (Chen et al., 2020) 63.41 69.77 73.53 60.84 68.27 73.39 54.98 56.97 62.52 51.41 69.44 77.73 59.78 68.52 76.54 59.02 67.26 73.20
BYOL (Grill et al., 2020) 71.70 73.83 76.45 57.16 67.44 71.64 48.73 61.63 70.82 41.99 74.40 77.17 60.88 74.42 78.75 54.20 71.92 74.69
BarlowTwins (Zbontar et al., 2021) 72.87 75.96 78.41 62.57 70.84 74.34 52.12 60.39 66.14 50.12 73.54 77.62 55.12 72.75 78.39 60.72 71.64 77.43
MoCo-v3 (Chen et al., 2021) 73.19 76.65 78.26 55.88 69.21 76.69 50.32 63.71 71.31 51.38 71.66 74.33 62.13 76.74 75.29 54.61 74.26 77.68
SimSiam (Chen & He, 2021) 73.15 72.70 75.63 62.52 69.31 76.38 55.16 62.91 71.31 49.30 69.47 75.92 58.35 72.89 75.31 58.25 68.61 77.41
TS-TCC (Eldele et al., 2021) 70.73 75.88 78.91 53.54 66.98 77.87 48.04 61.79 71.18 43.34 69.48 78.23 57.07 73.62 78.72 55.26 68.48 76.79
CLOCS (Kiyasseh et al., 2021) 68.94 73.36 76.31 57.94 72.55 76.24 51.97 57.79 72.65 47.19 71.88 76.31 59.59 77.78 77.49 54.38 71.93 76.13
ASTCL (Wang et al., 2023) 72.51 77.31 81.02 61.86 68.77 76.51 44.14 60.93 66.99 52.38 71.98 76.05 57.90 77.01 79.51 56.40 70.87 75.79
CRT (Zhang et al., 2023) 69.68 78.24 77.24 61.98 70.82 78.67 46.41 59.49 68.73 47.44 73.52 74.41 58.01 76.43 82.03 56.21 73.70 78.80
ST-MEM (Na et al., 2024) 61.12 66.87 71.36 54.12 57.86 63.59 55.71 59.99 66.07 51.12 65.44 74.85 56.69 63.32 70.39 59.77 66.87 71.36
MERL (Liu et al., 2024b) 82.39 86.27 88.67 64.90 80.56 84.72 58.26 72.43 79.65 53.33 82.88 88.34 70.33 85.32 90.57 66.60 82.74 87.95
D-BETA 83.15 88.36 90.11 77.74 82.92 85.15 70.10 78.91 83.98 86.61 92.83 96.71 85.46 91.35 94.92 80.04 87.36 90.71
@@
PhysioNet 2021. This contains ECG samples (500 Hz) ranging between 5 and 144 seconds. We process and fine-tune the subsets as described in (Oh et al., 2022) to validate the pre-trained ECG encoder in two downstream tasks: 1) 26-multi-label cardiac arrhythmia classification (Dx.); 2) patient identification (Id.), predicting patient ownership of ECG recordings.
PTB-XL. The PTB-XL dataset includes 21,837 ECG signals collected from 18,885 patients. Each sample has a 12-lead ECG recording sampled at 500 Hz over 10 seconds and corresponding cardiac labels. We follow (Liu et al., 2024b) to split this dataset, including four sub-groups (super, sub, form, and rhythm). We consider them as the four separated datasets and prepare each of them with the same train, val, and test set as in the original paper (Wagner et al., 2020).
CSN. This dataset consists of 23,026 ECG recordings sampled at 500 Hz for 10 seconds with 38 distinct labels, which also supports the evaluation in a classification task. We use 70%:10%:20% data split as processed in (Liu et al., 2024b).
CPSC2018. The dataset contains 6,877 standard 12-lead ECG recordings (500 Hz), which cover 9 distinct categories. Similarly, we also use the same data configuration following (Liu et al., 2024b).
CODE-test: This contains 827 12-lead ECG samples (400 Hz) at varying lengths covering 6 abnormalities, annotated by experienced residents and medical students. We resample the signals to 500 Hz and adjust the lengths to 10 seconds. We provide more detail about this dataset in Appendix A.1.
Table 3: Zero-shot performance (AUC in %) comparison across multiple datasets.
@@
D-BETA 76.2 75.9 66.1 88.6 80.1 76.3 77.1
Table 4: Zero-shot performance (AUC in %) under data distribution shift.
@@
MERL (Liu et al., 2024b) 0% 88.21 78.01 76.77 76.56 74.15 82.86
D-BETA 0% 72.09 79.11 77.12 82.91 76.24 80.10
Table 5: ECG interpretation comparison (AUC in %): Human experts vs. DNN (Ribeiro et al., 2020) vs. D-BETA .
Cardio Resident Emergency Resident Medical Student DNN D-BETA  (Zero-shot)
@@
Experimental Configurations.
To evaluate our model’s performance on downstream tasks, we conduct three experiments: 1) First, integrating a linear layer on top of the pre-trained ECG encoder and fine-tuning the entire model to test its efficacy in two tasks within the Physionet 2021 dataset: Dx. (by CinC score) and Id. (by % accuracy). We report the results with five cases of lead combinations, as presented in (Oh et al., 2022); 2) Second, we also implement a linear classifier but keep the ECG encoder frozen. This linear probing approach is applied at different training set sizes (1%, 10%, and 100%) to assess the macro AUC score (%) on the PTB-XL, CSN, and CPSC2018 test datasets, facilitating a comparison with our baseline (Liu et al., 2024b); 3) Finally, we investigate zero-shot classification (AUC) on PTB-XL, CSN, CPSC2018 and CODE-test datasets. Here, the texts used are obtained by passing the category names into GPT-4o to capture better medical context. The detailed configuration on each experiment is mentioned in Appendix A.1.
4.2 Experimental Results.
Full Fine-Tuning Evaluation.
As shown in Table 1, our method consistently outperforms previous approaches (Oh et al., 2022) in both examined tasks. In the classification task, our model achieves 85.7% accuracy with all 12 leads, significantly higher than the best baseline (W2V+CMSC+RLM), which is 73.2%. This number is even lower than our setting with only 1 lead usage (76.5%). Interestingly, the 3-lead combination yields the second-highest result, only 1.5% lower than using all leads, while the 2-lead and 6-lead combinations produce comparable results, both around 81.5%. This suggests that the selected leads (I, II, V2) capture sufficient information for accurate performance. A similar pattern emerges in the identification task, where our model achieves 41.1% accuracy with a single lead, 60.5% with 3 leads, and 65.4% with 12 leads, surpassing the best baseline by 7%.
Linear Probing Evaluation.
Zero-shot Evaluation.
We first compare our results and the best results of MERL in zero-shot settings across six datasets, as shown in Table 3. On average, our method achieves 77%, outperforming MERL by 2%. Notably, MERL performs impressively on the CPSC2018 dataset, while its results on the other five datasets are consistently lower than ours. Next, we extend the comparison of our method with MERL and other SSL baselines (Liu et al., 2024b) under data distribution shifts. Specifically, we compare linear probing (100% training size) of SSL methods with MERL’s and our zero-shot approach. In this setup, the source domain and target domain share some common categories. Details on this implementation can be found in Appendix A.1. As shown in Table 4, our results surpass MERL and other SSL methods, except when CPSC2018 is the target domain, which aligns with our previous observations. Finally, Table 5 shows that our zero-shot model outperforms three experienced cardiologists (over 3%) and also the in-domain model (Ribeiro et al., 2020), i.e., trained with millions of annotated ECG examples  [^1] Medical students outperform residents due to recent frequently focused training, aligning with prior work (Ribeiro et al., 2020).. We discuss more on zero-shot settings in Appendix A.3.
@@
To better understand how our method improves downstream performance, we visualize and compare the t-SNE embeddings generated by our ECG encoder on the CSN test set with those from MERL. For clearer visualization, we include only samples from unique categories and exclude categories with fewer than 50 samples. Figure 3 reveals that our embeddings show more well-defined and distinct clusters representing different ECG diagnoses.
Figure 3: T-SNE visualization on the CSN test set.
4.3 Ablation Studies.
We evaluate the impact of the key model components, the choice of language encoders, and varying the number of transformer layers in the ECG encoder for ablation studies. Here, we focus on three downstream tasks, including full fine-tuned diagnosis classification (results across five lead combinations), linear probing at 1% training size, and zero-shot classification (results across PTB-XL, CSN, and CPSC2018 datasets) using category name  [^2] We do not use GPT-4o for context enhancement, as our objective is to provide the core impact of our proposed components in the context of ablation studies..
Table 6: Effects of the model components: ① Flan-T5, ② ETS  Loss, ③ N3S.
@@
Effects of Key Components in D-BETA .
We systematically remove one component at a time from the default proposed model to assess the contribution of different model components, including Flan-T5, ETS , and N3S. Specifically, we start by eliminating the N3S and train the model with randomly selected negative samples. Subsequently, we take the ETS  loss away to assess its effectiveness in capturing rich representative embeddings in both encoders. Lastly, by replacing the Flan-T5 language encoder with a standard Bert-base architecture (Devlin, 2018), we consider this as the baseline model. Table 6 demonstrates the results of this experiment. It can be seen that ETS  significantly enhances performance, showing an improvement of approximately 15% in both full fine-tuning and linear probing settings over the baseline model. Meanwhile, adding N3S improves zero-shot classification by 2%, and introducing Flan-T5 enhances performance in linear probing by 4% compared to the baseline. These results underscore the effectiveness of each component in optimizing the model’s performance.
Text Encoders.
Table 7: Effects of different text encoders.
Text encoder Full fine-tune Linear probing Zero-shot
@@
Number of ECG Transformer Layers.
As part of our ablation study, we explore the impact of varying the number of transformer layers (1, 4, 8) in the ECG encoder. As shown in Table 8, increasing the number of layers significantly improves performance. Specifically, the 1-layer model performs 11% worse than the 8-layer model (proposed) in full fine-tuning and 13% worse in linear probing. For zero-shot, the 8-layer model still delivers superior results, with 2% and 3% higher performance than the 4-layer and 1-layer models, respectively. Although these differences are smaller than in full fine-tuning, they still highlight the ECG encoder design’s impact on improving performance. It is also worth noting that in this same zero-shot setting (without GPT-4o’s support), even our 1-layer case outperformed MERL by approximately 8%, achieving approximately 70% compared to MERL’s 62%.
Table 8: Effects of number of transformer layers in ECG encoder.
@@ 5 Conclusion
We propose D-BETA , a novel contrastive masked transformer-based architecture to pre-train ECG signals and corresponding texts. Our approach is generative self-supervised learning, enhanced with ETS  loss, and nearest-neighbor negative sampling to support contrastive aspects. Experimental results demonstrate that our method outperforms previous approaches across multiple datasets and on a wide range of downstream tasks with over 100 cardiac conditions. D-BETA  shows promise in advancing ECG-based diagnostic models, paving the way for more accurate, efficient, and personalized cardiac care.
Impact Statement
Our D-BETA  framework advances ECG diagnosis by leveraging multi-modal self-supervised learning, delivering robust benchmarking results across various downstream tasks. We hope D-BETA  provides valuable modeling insights for ECG diagnosis and contributes to the healthcare community focused on cardiovascular diseases. We also publicly release the pre-trained models and code, empowering future research and extending ECG analysis to broader clinical applications such as ECG question answering and retrieval-based reporting.
Acknowledgments
This research was supported by the Google South Asia & Southeast Asia research award.
@@ Appendix A Appendix
We first visualize representative examples of ECG-text pairs from the MIMIC IV ECG dataset (Gow et al., 2023), as shown in Figure 4. We also indicate the top 30 common unique reports (before merging) in Figure 5. Prominent terms such as ”abnormal ecg”, ”normal ecg”, ”atrial fibrillation”, and ”sinus tachycardia” indicate common diagnoses, which suggests prevalent cardiovascular conditions and typical annotations within this dataset.
@@
Figure 5: WordCloud visualization on the top 30 common unique reports from the MIMIC IV ECG dataset.
Next, we provide more details on data configurations in Table 9, including data split, number of classes, metrics, and the corresponding tasks with the given downstream datasets. Particularly, the CODE-test dataset is from the work (Ribeiro et al., 2020) with their model being trained on a set of over 2 million ECG records from 1,676,384 different patients in 811 counties. We show our model’s effectiveness by evaluating our performance on the same released test set of 827 samples, but in a zero-shot manner. The samples are originally sampled at 400 Hz, with durations of either 10 seconds or 7 seconds. Therefore, we resampled to 500 Hz and adjusted by truncating or padding with zeros as needed to get 10-second samples. For the gold standard (ground truth), two expert cardiologists provided their diagnoses. If they agree with each other, their consensus becomes the gold standard. In cases of disagreement, a third specialist reviews their diagnoses and determines the final decision.
@@
Table 10: Details on training configurations on the fine-tuned datasets. For the optimizer, we keep using Adam in all experiments.
@@
Finally, for the distribution shift experiment, we follow the SCP-codes (classes) matching settings in (Liu et al., 2024b), as shown in Table 11. This is to support three dataset matches (PTBXL-Super and CPSC2018), (PTBXL-Super and CSN), and (CPSC2018 and CSN). It is worth noting that the None value indicates the target dataset does not have a matching label for given labels in the source dataset.
@@
Why Not Use ETM Only in Zero-shot Learning.
As mentioned in the Method section, ETM functions as a contrastive learning technique in the masked auto-encoder architecture. However, it heavily relies on binary classification tasks with explicit ECG-text pairs to learn cross-modal correspondences. It is not designed for zero-shot learning which strongly requires the model to generalize to unseen tasks or classes without the need for such supervised pairings or fused information during training. This motivates us to use ETS , boosting the model’s zero-shot ability.
@@
In medical datasets, particularly the MIMIC-IV ECG dataset (Gow et al., 2023), a significant amount of duplicate or highly similar text samples: among nearly 800,000 records, only approximately 180,000 are unique. For instance, over 100,000 samples share an identical text report, which is ”sinus rhythm normal ecg”. Randomly selecting negative samples for contrastive loss training is not a suitable approach in this scenario. Therefore, we propose using the N3S technique to more effectively differentiate between similar and dissimilar samples, improving contrastive learning by selecting more meaningful negatives. Notably, during training, we observe that the ETM accuracy without N3S stagnates around 75%, while with N3S, it exceeds 96%, demonstrating the significant impact of this approach.
A.3 Enhancing Zero-shot Performance with LLMs.
Limiations in MERL’s Approach.
@@
In zero-shot learning, models typically rely on category names alone to make predictions. However, by incorporating Large Language Models (LLMs), we can enhance the context by generating richer, clinically relevant descriptions of the categories, as discussed in MERL (Liu et al., 2024b). However, we observe two main drawbacks in their enhanced text reports, as shown in Figure 6: 1) MERL’s performance heavily depends on their sub-types and attributes searching prompt and additional database. This leads to a limitation when testing detailed analysis with labels that are different sub-types themselves. Moreover, this also raises suspicion about the performance when new tasks require labels that are not able to search sub-types and attributes in the database; 2) Following that point, MERL’s enhanced texts might be uncontrollable to the outputs where the LLMs provide wrong sub-types or unnecessary context. For example, ”Atrial Fibrillation” is already in ”AFIB” type but shown to misleadingly be in ”AF”-”Atrial Flutter” in their settings.
LLMs in D-BETA .
We address these above points using a straightforward prompt strategy with explicit instructions. Specifically, we employ a prompt: ”You are an experienced cardiologist. For a given clinical term such as ’normal ECG’, your job is to describe each term clinically and apply your medical domain knowledge to include other relevant explanations that will help a text encoder like Flan-T5 fully understand medical concepts. Do not include any recommendations in the description.” This makes the LLM generate clinically accurate and more focused explainable descriptions, enhancing the text encoding without introducing irrelevant or redundant information. For example, with the code ”AFIB”, our prompt on GPT-4o can output: ”Atrial Fibrillation (AFIB). Irregular and often rapid heart rate due to uncoordinated atrial activity.”.
Additional Experiments.
Here, we present additional experiments to highlight the effectiveness of ETM loss and masking modeling techniques (e.g., MLM, MEM). Specifically, we perform zero-shot classification with GPT-4o support (reported in AUC (%)) on four datasets: PTBXL-Super, PTBXL-Form, CSN, and CODE-Test.
Table 12: Impact of ETM. Results report zero-shot classification in AUC (%).
PTBXL-Super PTBXL-Form CSN CODE-Test
w/o ETM 73.2 65.8 76.6 96.2
w ETM 76.2 66.1 76.3 96.8
As indicated in Table 12, the impact of ETM is demonstrated where, removing ETM slightly decreases performance across most datasets, particularly in PTBXL-Super (76.2 to 73.2). However, the effect on CSN is minimal, suggesting dataset-specific sensitivity to ETM.
Table 13: Impact of MLM and MEM. Results report zero-shot classification in AUC (%).
PTBXL-Super PTBXL-Form CSN CODE-Test
w/o MLM + MEM 70.3 67.4 74.5 94.6
w MLM + MEM 76.2 66.1 76.3 96.8
Finally, we can observe that incorporating MLM and MEM noticeably improves performance across all evaluated datasets in Table 13. Especially, gains are observed in PTBXL-Super (+5.9%), and CODE-Test (+2.2%), demonstrating that the reconstruction tasks also play an important role in enhancing the model’s ability for better performance.